OpenStudy (anonymous):

Tom Petty was born 6 years before Tom Hanks. Tom Cruise was born 6 years after Tom Hanks. Let x represent the age of Tom Hanks. Write algebraic expressions to represent the ages of the other two celebrities.

OpenStudy (anonymous):

tom petty = x+6 tom cruise = x - 6 tom hanks = x

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

A married couple needed to purchase 28 presents for friends and relatives on their holiday gift list. If the husband purchased g presents, write an algebraic expression that represents the number of presents that the wife needs to buy.

OpenStudy (anonymous):

y=28-x where x=husband bought and y=wife